# Utilizing Jailbreak Probability to Attack and Safeguard Multimodal LLMs

### v1

@@ Abstract
Recently, Multimodal Large Language Models (MLLMs) have demonstrated their superior ability in understanding multimodal contents. However, they remain vulnerable to jailbreak attacks, which exploit weaknesses in their safety alignment to generate harmful responses. Previous studies categorize jailbreaks as successful or failed based on whether responses contain malicious content. However, given the stochastic nature of MLLM responses, this binary classification of an input’s ability to jailbreak MLLMs is inappropriate. Derived from this viewpoint, we introduce jailbreak probability to quantify the jailbreak potential of an input, which represents the likelihood that MLLMs generated a malicious response when prompted with this input. We approximate this probability through multiple queries to MLLMs. After modeling the relationship between input hidden states and their corresponding jailbreak probability using Jailbreak Probability Prediction Network (JPPN), we use continuous jailbreak probability for optimization. Specifically, we propose Jailbreak-Probability-based Attack (JPA) that optimizes adversarial perturbations on inputs to maximize jailbreak probability. To counteract attacks, we also propose two defensive methods: Jailbreak-Probability-based Finetuning (JPF) and Jailbreak-Probability-based Defensive Noise (JPDN), which minimizes jailbreak probability in the MLLM parameters and input space, respectively. Extensive experiments show that (1) JPA yields improvements (up to 28.38%) under both white and black box settings compared to previous methods with small perturbation bounds and few iterations. (2) JPF and JPDN significantly reduce jailbreaks by at most over 60%. Both of the above results demonstrate the significance of introducing jailbreak probability to make nuanced distinctions among input jailbreak abilities. Disclaimer: this paper may include harmful content.
1 Introduction
Multimodal Large Language Models (MLLMs) exhibit excellent comprehension of multimodal contents with billions of parameters and massive training data, achieving high performance across multiple multimodal tasks, such as visual question answering [19, 7], visual grounding [49, 26], and image captioning [30, 52], etc. However, these MLLMs are vulnerable to jailbreak attacks, which are techniques designed to compromise model safety alignment and induce harmful or undesirable outputs, such as hate speech or discriminatory comments. Although jailbreak attacks on LLMs have been widely studied [54, 53, 20, 46], research on these vulnerabilities in MLLMs remains relatively limited [34, 21, 11, 3].
@@ 1 Introduction
Figure 2: Illustration of JPA, JPDN and JPF. ALL of three methods first extract hidden states from victim MLLM. Then (1) JPA maximizes jailbreak probability and update input image. (2) JPDN minimizes jailbreak probability and update input image. (3) JPF minimizes jailbreak probability and update victim MLLM.
Leveraging this metric, we introduce the Jailbreak Probability Prediction Network (JPPN) to model the relationship between jailbreak probability and hidden states of inputs as illustrated in Figure 1. This enables the optimization of input to control their jailbreak potential. Specifically, we propose Jailbreak-Probability-based Attack (JPA), which updates the input image through optimizing hidden states to maximize jailbreak probability. To mitigate jailbreak, we also propose two defense methods to minimize jailbreak probability in the MLLM parameters and input space. Similar to adversarial training [27], Jailbreak-Probability-based Finetuning (JPF) fine-tunes MLLM to reduce jailbreak probability in adversarial images. Jailbreak-Probability-based Defensive Noise (JPDN) modifies the input image to introduce defensive noises, which also reduce jailbreak probability. The effectiveness of Jailbreak-Probability-based methods comes from the nuanced jailbreak probability labels of inputs, which enable JPPN to learn the key features for both conducting and defending against jailbreaks in feature space. The illustration of Jailbreak-Probability-based methods is shown in Figure 2.
We conduct extensive experiments to demonstrate the advantages of JPA. The results show that JPA outperforms previous methods in attack success rate (ASR) on five MLLMs under both white-box and black-box settings, using only 1,000 epochs and a perturbation bound of 16/255. Additional experiments show that JPA maintains strong jailbreak performance even with a small perturbation bound of 4/255. Visualization of tokens from hidden states further confirms that JPA successfully modifies the hidden states. Moreover, experiments evaluating the defense performance of JPDN and JPF show that both methods reduce the ASR by at most over 60%. These findings demonstrate the effectiveness of leveraging jailbreak probability for both attacking and safeguarding MLLMs.
@@
2 Related Work
2.1 Multimodal Large Language Models
@@
Recently, MLLMs have shown their great performance on perception and comprehension across various multimodal tasks, and become a significant area of research [44]. MLLMs usually contain a visual encoder, a connector for projecting visual embeddings to text representation space and an LLM for generating responses. MiniGPT-4 [52], InstructBLIP [7], Qwen-VL [1] ,InternLM-XComposer-VL [48] and DeepSeek-VL [25] are typical open-source MLLMs and are employed in our study.
2.2 Jailbreak Attack
@@
In the context of LLMs, jailbreak attacks refer to the techniques that use adversarial examples to bypass safety alignment and mislead LLMs into generating harmful or unwanted responses [43]. There are many previous methods conducting jailbreak attacks on LLMs using different mechanisms, such as GCG [54], AutoDAN [53], DeepInception [20] and CipherChat [46]. Similar to LLMs, MLLMs are also vulnerable to jailbreak attacks [4, 34, 10, 21, 11, 3, 45].Jailbreak attacks on MLLMs can be roughly divided into two categories [23]: adversarial attacks and visual prompt injections. Adversarial attacks usually set a loss function and use the gradients of loss w.r.t input to generate adversarial images [4, 34, 10], while visual prompt injections add malicious texts to the input image and leverage the OCR ability of MLLMs to conduct jailbreak attack [11]. There are also hybrid methods utilizing both of the above mechanisms [21, 37]. Besides, leveraging text-to-image models (e.g. Stable Diffusion [36]) to generate query-relevant images is also a widely adopted technique to jailbreak MLLMs [21, 22]. Although these methods achieve satisfactory results, they still have unresolved drawbacks, such as high time costs when generating adversarial examples and poor invisibility, thus need improvements.
2.3 Jailbreak Defense
@@
As the development and broad applications of (M)LLMs continue to progress, the safety of them have become a significant concern [42, 15]. Alignment techniques are proposed to fulfill the goal of letting (M)LLMs follow users’ instructions and be helpful, truthful and harmless [31], and protect them from jailbreak attacks. Reinforcement Learning from Human Feedback (RLHF) is a commonly used method for safety alignment and has demonstrated its effectiveness [31, 2, 6, 16, 50, 32, 14]. Besides, there are other defenses for MLLMs such as employing auxiliary modules to detect harmful outputs[33], utilizing prompts to enhance safety [13, 39], or distinguishing the differences of distributions between benign and harmful inputs [41].
3 Method
3.1 Problem Formulation
@@
We consider a single-turn conversation between an attacker and an MLLM, where the attacker aims to execute a jailbreak attack by bypassing the MLLM’s safety alignment and prompting it to generate harmful responses. We assume a white-box setting, in which the attacker has full access to the MLLM. We denote the MLLM as $M$, which takes an image-text pair $X=(x_{img},x_{txt})$ from dataset $D$ as its input and gives a textual response $y=M(X)$ in an autoregressive manner. A safety criterion $C$ (e.g. manual inspection or a Judge LLM) assigns a binary safety score $s\in\{0,1\}$ to indicate whether $y$ is harmful, where $s=1$ denotes harm and $s=0$ denotes harmlessness. The attempt of jailbreaking $M$ is formulated as follows:
@@
where $X^{\prime}=(x^{\prime}_{img},x^{\prime}_{txt})$ is an adversarial example. $x^{\prime}_{img}$ is represented by the following equation:
$x^{\prime}_{img}=x_{img}+\delta,\quad\text{s.t.}\quad\|\delta\|_{\infty}\leq\epsilon$ (2)
where $\delta$ is the added perturbation and $\epsilon$ is used to restrict the magnitude of $\delta$. Since we only consider altering the image modality in our study, we have $x^{\prime}_{txt}=x_{txt}$.
3.2 Jailbreak Probability
Previous adversarial attacks on discriminative tasks (e.g. image classification [12, 28, 9] and image-text retrieval [47, 24, 40]) have definite results whether an attack is successful. However, jailbreak attacks target generative tasks, where the MLLM responses to one specific input are stochastic, making it both inappropriate and impossible to determine whether the input is able to jailbreak the MLLM. To support this viewpoint, we generate 40 responses for every sample in Zhou’s dataset [51] (see Section 4.1 for details) using MiniGPT-4 and count how many times each sample succeeds in jailbreak, as depicted in Figure 3.
@@ 3.2 Jailbreak Probability
The jailbreak probability $P^{M}_{X}$ is defined as the probability that $M$ generates a jailbreak response when prompted with $X$:
@@
where $N$ represents the total number of responses sampled from $M$.
However, determining the exact value of $P^{M}_{X}$ requires infinite samples. An alternative is to approximate it by generating responses for $n$ times and calculating the proportion of successful jailbreaks. This empirical estimate, referred to as the approximated jailbreak probability and denoted as $\hat{p}_{X}^{M}$ ($n$ is left out for simplicity), provides a practical measure of $P^{M}_{X}$:
$\hat{p}_{X}^{M}=\frac{1}{n}\sum_{i=1}^{n}C(M(X))$ (4)
Theoretically, $\hat{p}_{X}^{M}$ will gradually converge to $P^{M}_{X}$ as $n$ grows larger.
3.3 Jailbreak Probability Prediction Network
Given the role of jailbreak probability in fine-grained measurement of jailbreak ability, modeling its relationship with input hidden states enables a deeper understanding of MLLM vulnerabilities. We propose the Jailbreak Probability Prediction Network (JPPN), which takes the hidden states of $X$ in $M$ as the input to predict its jailbreak probability, to model the inherent correlation between $\hat{p}_{X}^{M}$ and $X$. Compared to the representation from foundational models (e.g. CLIP [35]), these hidden states capture the internal understanding of $X$ from $M$. Since hidden states are already highly encoded by $M$, we simply use an MLP to implement JPPN and link it to each Transformer block. We denote JPPN as $J$, with its output $j_{X}^{M}\in[0,1]$ representing the predicted jailbreak probability of $X$ on $M$:
@@ 3.3 Jailbreak Probability Prediction Network
where $HS_{M}(X)$ is the hidden states obtained by passing $X$ through $M$. We train the JPPN by minimizing the following MSE loss:
$\mathcal{L}_{train}=\frac{1}{\left|D\right|}\sum_{X\in D}\mathcal{L}_{MSE}(j_{X}^{M},\hat{p}_{X}^{M})$ (6)
The training procedure is illustrated in Figure 1. After training, it is important to verify the accuracy of JPPN to ensure its capability of distinguishing different levels of jailbreak probability. Given the inherent randomness in the value of $\hat{p}_{X}^{M}$, we introduce a flexible metric for evaluating accuracy, which is referred to as $\tau$-thresholded accuracy ($acc_{\tau}$). $acc_{\tau}$ is calculated with an indicator function:
$acc_{\tau}=\frac{1}{\left|D\right|}\sum_{X\in D}\mathbf{1}(X),\,\,if\left|j_{X}^{M}-\hat{p}_{X}^{M}\right|\leq\tau$ (7)
The results in Figure 4 demonstrate that JPPN effectively predicts jailbreak probability.
3.4 Jailbreak Probability for Attack
Once the correlation of $\hat{p}_{X}^{M}$ and the input $X$ is effectively encoded, it can be leveraged to guide inputs towards achieving higher jailbreak probability. With a trained JPPN with high accuracy, we propose the Jailbreak-Probability-based Attack (JPA) which generates the adversarial example $X^{\prime}$ by minimizing the following loss:
$\mathcal{L}_{JPA}=\mathcal{L}_{MSE}(j_{X}^{M},1)$ (8)
where $1$ is the target and represents 100% probability of $X$ jailbreaking $M$. The adversarial image $x^{\prime}_{img}$ is generated iteratively using BIM [18] via the following equation:
@@ 3.4 Jailbreak Probability for Attack
s.t. $\quad\|x^{\prime}_{img}-x_{img}\|_{\infty}\leq\epsilon$ (10)
where $x^{\prime}_{img}$ is initialized as $x_{img}$, $\alpha$ is the step size, $\epsilon$ is the perturbation bound and $sign$ is sign function. After several iterations of updates, the hidden states of $X^{\prime}$ are located close to those of high jailbreak probability, thereby enhancing the likelihood that $X^{\prime}$ will successfully jailbreak $M$. The illustration of JPA is shown in Figure 2 .
The effectiveness of JPA comes from the nuanced jailbreak probability labels of malicious inputs, which enable JPPN to learn what is crucial for achieving successful jailbreak in feature space. The essence of jailbreak probability is a kind of additional information about inputs, which brings knowledge about the inherent vulnerabilities of the victim MLLM. Furthermore, compared to methods that align the outputs of models with harmful responses [34, 45, 21], JPA utilizes a small MLP to perform optimization on a simple 1-digit regression task, which is fast and much easier to converge. As shown in Section 4.3, JPA can achieve high attack success rates with a small perturbation bound and few epochs. The algorithm for JPA is presented in Algorithm 1.
Algorithm 1 Jailbreak-Probability-based Attack (JPA)
0:  image-text pair input $X=(x_{img},x_{txt})$, white-box MLLM $M$, pre-trained JPPNs $J$, hidden states extractor $HS_{M}$, learning rate $\alpha$, sign function $sign$, value clipping operation $Clip_{\epsilon}$, iterations $I$
0:  adversarial image $x^{\prime}_{img}$
1:  $x^{\prime}_{img}=x_{img}$
2:  for $t=1$ to $I$ do
3:     $j_{X^{\prime}}^{M}=J(HS_{M}(x^{\prime}_{img},x_{txt}))$
4:     $\mathcal{L}_{JPA}=\mathcal{L}_{MSE}(j_{X^{\prime}}^{M},1)$
6:  end for
7:  return  $x^{\prime}_{img}$
3.5 Jailbreak Probability for Defense
Given the more effective guidance provided by jailbreak probability, instead of minimizing the difference between $j_{X}^{M}$ and $1$ in JPA (Eq.8), we can drive $j_{X}^{M}$ closer to 00 to defend against potential jailbreaks:
@@ 3.5 Jailbreak Probability for Defense
where 00 represents 0% probability of $X$ jailbreaking $M$. By minimizing Eq. 11 on MLLM parameters or the input space, we propose two defense methods: Jailbreak-Probability-based Finetuning (JPF) and Jailbreak-Probability-based Defensive Noise (JPDN). JPF finetunes MLLMs via:
@@
where $\theta$ are the parameters of $M$ and $\beta$ represents the learning rate. For JPDN, images are updated in the same way as JPA, where the perturbation $\delta$ is applied to mitigate the jailbreak effect. We illustrate JPF and JPDN in Figure 2. In practice, JPF can be leveraged during safety alignment, while JPDN can be served as input image pre-processsing, to enhance the safety of models.
4 Experiments
4.1 Experimental Setup
@@
Datasets. Zhou et al. [51] introduce a malicious question dataset which selectively consists of three datasets, including AdvBench [54], StrongREJECT dataset [38] and JailbreakBench [5]. We refer to this dataset as Zhou’s dataset. Zhou’s dataset is randomly split into train set and test set with ratios of 0.8 and 0.2. For all texts in the Zhou’s dataset, we combine them with a photo of a panda to form image-text pairs, which is the same as VAP [34] and BAP [45]. We also use MM-SafetyBench [22] in our experiments.
@@
Models. We adopt MiniGPT-4 (Vicuna 13B) [52], InstructBlip (Vicuna 13B) [7], Qwen-VL-Chat (Qwen-7B) [1] and InternLM-XComposer-VL (InternLM-Chat-7B) [48], DeepSeek-VL-1.3B-Chat (DeepSeek-LLM-1.3B) [25] in our experiments.
@@
Criterion. We use OpenAI’s GPT-4 as the judging model, which gives judgments of ”safe” or ”unsafe” to the outputs of models. See Supp. B for the details of the judging prompt. In order to reduce randomness, we generate answers for 5 times in all experiments.
4.2 Training of JPPNs
@@
Training Settings. We choose MiniGPT-4 as the MLLM for implementing JPPN. JPPN is implemented as a three-layer MLP. According to the dimensions of MiniGPT-4’s hidden states, JPPN’s input dimension is set to 5120. Each Transformer block is connected to a distinct JPPN to process its respective hidden states. Details on the complete JPPN architecture can be found in Supp. A. We train JPPNs on Zhou’s train set. To label malicious questions, we generate multiple MiniGPT-4 responses and calculate the approximated jailbreak probability using Equation 4 based on GPT-4. Since jailbreak probability is derived from the ratio of successful jailbreak responses, its stability depends on the number of responses. To assess the impact of the response number on training, we experiment with 5, 20, and 40 responses per input, training JPPNs-5r, JPPNs-20r, and JPPNs-40r, respectively, over 150 epochs. We use Adam [17] with an initial learning rate of 0.001 as the optimizer, and reduce the learning rate to 20% every 50 epochs. Each JPPN performs optimization independently by minimizing Equation 8, allowing precise accuracy evaluation for individual JPPNs and enabling flexible combinations.

### v4

# Probabilistic Modeling of Jailbreak on Multimodal LLMs: From Quantification to Application

@@ Abstract
Recently, Multimodal Large Language Models (MLLMs) have demonstrated their superior ability in understanding multimodal content. However, they remain vulnerable to jailbreak attacks, which exploit weaknesses in their safety alignment to generate harmful responses. Previous studies categorize jailbreaks as successful or failed based on whether responses contain malicious content. However, given the stochastic nature of MLLM responses, this binary classification of an input’s ability to jailbreak MLLMs is inappropriate. Derived from this viewpoint, we introduce jailbreak probability to quantify the jailbreak potential of an input, which represents the likelihood that MLLMs generated a malicious response when prompted with this input. We approximate this probability through multiple queries to MLLMs. After modeling the relationship between input hidden states and their corresponding jailbreak probability using Jailbreak Probability Prediction Network (JPPN), we use continuous jailbreak probability for optimization. Specifically, we propose Jailbreak-Probability-based Attack (JPA) that optimizes adversarial perturbations on input image to maximize jailbreak probability, and further enhance it as Multimodal JPA (MJPA) by including monotonic text rephrasing. To counteract attacks, we also propose Jailbreak-Probability-based Finetuning (JPF), which minimizes jailbreak probability through MLLM parameter updates. Extensive experiments show that (1) (M)JPA yields significant improvements when attacking a wide range of models under both white and black box settings. (2) JPF vastly reduces jailbreaks by at most over 60%. Both of the above results demonstrate the significance of introducing jailbreak probability to make nuanced distinctions among input jailbreak abilities.
Disclaimer: this paper may include harmful content.
Introduction
Multimodal Large Language Models (MLLMs) exhibit excellent comprehension of multimodal contents with billions of parameters and massive training data, achieving high performance across multiple multimodal tasks, such as visual question answering (Li et al. 2023a; Dai et al. 2023b), visual grounding (Zhao et al. 2023; Ma et al. 2024), and image captioning (OpenAI 2024; Zhu et al. 2023a), etc. However, these MLLMs are vulnerable to jailbreak attacks, which are techniques designed to compromise model safety alignment and induce harmful or undesirable outputs, such as hate speech or discriminatory comments. Although jailbreak attacks on LLMs have been widely studied (Zou et al. 2023; Zhu et al. 2023b; Li et al. 2023b; Yuan et al. 2024), research on these vulnerabilities in MLLMs remains relatively limited (Qi et al. 2024; Li et al. 2024; Gong et al. 2023; Bailey et al. 2024).
@@ Introduction
Figure 2: Illustration of (M)JPA and JPF. All of these methods first extract hidden states from victim MLLM. Then (1) JPA maximizes jailbreak probability and update input image. (2) MJPA further conducts monotonic text refinement on the results of JPA. (3) JPF minimizes jailbreak probability and update victim MLLM.
Leveraging this metric, we introduce the Jailbreak Probability Prediction Network (JPPN) to model the relationship between jailbreak probability and hidden states of inputs as illustrated in Fig. 1. This enables the optimization of input to control their jailbreak potential. Specifically, we propose Jailbreak-Probability-based Attack (JPA), which updates the input image through optimizing hidden states to maximize jailbreak probability. To enhance JPA, we further incorporate a monotonic text rephrasing strategy that also boosts the jailbreak probability, thereby forming the Multimodal JPA (MJPA). To mitigate jailbreak, we also propose Jailbreak-Probability-based Finetuning (JPF) that fine-tunes MLLM to reduce jailbreak probability. The effectiveness of Jailbreak-Probability-based methods comes from the nuanced jailbreak probability labels of inputs, which enable JPPN to learn the key features for both conducting and defending against jailbreaks in feature space. The illustration of (M)JPA and JPF is shown in Fig. 2.
We conduct extensive experiments to demonstrate the advantages of (M)JPA. The results show that MJPA outperforms previous methods in attack success rate (ASR) on both open-source and commercial models, and that JPA is effective when applied to various forms of inputs. Additional experiments show that JPA maintains strong jailbreak performance even with a small perturbation bound of 4/255. Visualization of tokens from hidden states further confirms that JPA successfully modifies the hidden states. Moreover, experiments evaluating the defense performance of JPF show that JPF reduces the ASR by at most over 60%. These findings demonstrate the effectiveness of leveraging jailbreak probability for both attacking and safeguarding MLLMs.
@@
Related Work
Multimodal Large Language Models
@@
Recently, MLLMs have shown their great performance on perception and comprehension across various multimodal tasks, and become a significant area of research (Yin et al. 2023). MLLMs usually contain a visual encoder, a connector for projecting visual embeddings to text representation space and an LLM for generating responses. MiniGPT-4 (Zhu et al. 2023a), InstructBLIP (Dai et al. 2023b), Qwen-VL (Bai et al. 2023) ,InternLM-XComposer-VL (Zhang et al. 2023) and DeepSeek-VL (Lu et al. 2024) are typical open-source MLLMs and are employed in our study.
Jailbreak Attack
@@
In the context of LLMs, jailbreak attacks refer to the techniques that use adversarial examples to bypass safety alignment and mislead LLMs into generating harmful or unwanted responses (Yi et al. 2024). There are many previous methods conducting jailbreak attacks on LLMs using different mechanisms, such as GCG (Zou et al. 2023), AutoDAN (Zhu et al. 2023b), DeepInception (Li et al. 2023b) and CipherChat (Yuan et al. 2024). Similar to LLMs, MLLMs are also vulnerable to jailbreak attacks (Carlini et al. 2024; Qi et al. 2024; Dong et al. 2023; Li et al. 2024; Gong et al. 2023; Bailey et al. 2024; Ying et al. 2024).Jailbreak attacks on MLLMs can be roughly divided into two categories (Liu et al. 2024b): adversarial attacks and visual prompt injections. Adversarial attacks usually set a loss function and use the gradients of loss w.r.t input to generate adversarial images (Carlini et al. 2024; Qi et al. 2024; Dong et al. 2023), while visual prompt injections add malicious texts to the input image and leverage the OCR ability of MLLMs to conduct jailbreak attack (Gong et al. 2023). There are also hybrid methods utilizing both of the above mechanisms (Li et al. 2024; Shayegani, Dong, and Abu-Ghazaleh 2023). Besides, leveraging text-to-image models (e.g. Stable Diffusion (Rombach et al. 2022)) to generate query-relevant images is also a widely adopted technique to jailbreak MLLMs (Li et al. 2024; Liu et al. 2024a). Although these methods achieve satisfactory results, they still have unresolved drawbacks, such as high time costs when generating adversarial examples and poor invisibility, thus need improvements.
Jailbreak Defense
@@
As the development of (M)LLMs continue to progress, the safety of them have become a significant concern (Yao et al. 2024; Huang et al. 2023). Alignment techniques are proposed to fulfill the goal of letting (M)LLMs follow users’ instructions and be helpful, truthful and harmless (Ouyang et al. 2022), and protect them from jailbreak attacks. Reinforcement Learning from Human Feedback (RLHF) is a commonly used method for safety alignment and has demonstrated its effectiveness (Ouyang et al. 2022; Bai et al. 2022; Dai et al. 2023a; Ji et al. 2024; Zhou et al. 2024a; Pi et al. 2024a; Helff et al. 2024). Besides, there are other defenses for MLLMs such as employing auxiliary modules to detect harmful outputs(Pi et al. 2024b), utilizing prompts to enhance safety (Gou et al. 2024; Wang et al. 2024), or distinguishing the differences of distributions between benign and harmful inputs (Xu et al. 2024b).
Method
Problem Formulation
@@
We consider a conversation between an attacker and an MLLM, where the attacker aims to execute jailbreak by bypassing the MLLM’s safety alignment and inducing it to generate harmful responses. We assume a white-box setting, where the attacker has full access to the MLLM. We denote the MLLM as $M$, which takes an image-text pair $X=(x_{img},x_{txt})$ from dataset $D$ as input and gives a response $y=M(X)$ in an autoregressive manner. A safety criterion $C$ (e.g. manual inspection or a Judge LLM) gives a binary safety score $s\in\{0,1\}$ to indicate whether $y$ is harmful, where $s=1$ denotes harm and $s=0$ denotes harmlessness. The attempt of jailbreak is formulated as follows:
@@
where $X^{\prime}=(x^{\prime}_{img},x^{\prime}_{txt})$ is a crafted jailbreak example.
Jailbreak Probability
Previous adversarial attacks on discriminative tasks (e.g. image classification (Goodfellow 2014; Madry et al. 2018; Dong et al. 2018) and image-text retrieval (Zhang, Yi, and Sang 2022; Lu et al. 2023; Xu et al. 2024a)) have definite results on whether an attack is successful. However, jailbreak attacks target generative tasks, where the MLLM responses to one specific input are stochastic, making it both inappropriate and impossible to determine whether the input is able to jailbreak the MLLM. To support this viewpoint, we generate 40 responses for every sample in Zhou’s dataset (Zhou et al. 2024b) (see Section Experiments for details) using MiniGPT-4 and count how many times each sample succeeds in jailbreak, as depicted in Fig. 3.
@@
The jailbreak probability $P^{M}_{X}$ is defined as the probability that $M$ generates a harmful response with $X$:
@@
where $N$ is the total number of responses sampled from $M$.
However, determining the exact value of $P^{M}_{X}$ requires infinite responses. An alternative is to approximate it by generating responses for $n$ times and calculating the proportion of successful jailbreaks. This empirical estimate, referred to as the approximated jailbreak probability and denoted as $\hat{p}_{X}^{M}$ ($n$ is left out for simplicity), provides a practical measure of $P^{M}_{X}$:
$\hat{p}_{X}^{M}=\frac{1}{n}\sum_{i=1}^{n}C(M(X)).$ (3)
We can find that $\mathbb{E}[\hat{p}_{X}^{M}]=P^{M}_{X}$, which means that $\hat{p}_{X}^{M}$ is an unbiased estimator of $P^{M}_{X}$, and that the variance of $\hat{p}_{X}^{M}$ is $\frac{1}{n}(P^{M}_{X}(1-P^{M}_{X}))$. The above properties indicate that $\hat{p}_{X}^{M}$ will gradually converge to $P^{M}_{X}$ as $n$ grows larger.
Jailbreak Probability Prediction Network
Given the role of jailbreak probability in fine-grained measurement of jailbreak ability, modeling its relationship with input hidden states enables a deeper understanding of MLLM vulnerabilities. We propose the Jailbreak Probability Prediction Network (JPPN), which takes the hidden states of $X$ in $M$ as the input to predict its jailbreak probability, to model the inherent correlation between $\hat{p}_{X}^{M}$ and $X$. Compared to the representation from foundational models (e.g. CLIP (Radford et al. 2021)), these hidden states capture the internal understanding of $X$ from $M$. Since hidden states are already highly encoded by $M$, we simply use an MLP to implement JPPN and link it to each Transformer block. We denote JPPN as $J$, with its output $j_{X}^{M}\in[0,1]$ representing the predicted jailbreak probability of $X$ on $M$:
@@
where $HS_{M}(X)$ is the hidden states obtained by passing $X$ through $M$. We minimize the following loss to train JPPN:
$\mathcal{L}_{train}=\frac{1}{\left|D\right|}\sum_{X\in D}\mathcal{L}_{MSE}(j_{X}^{M},\hat{p}_{X}^{M}).$ (5)
The training procedure is illustrated in Fig. 1.
Theorem 1 (Generalization Error Bound for JPPN).
Let $P^{M}_{X}$ be the true jailbreak probability for an input $X$, and $j^{M}_{X}$ be its estimate by JPPN. For $\delta\in(0,1)$, with probability at least $1-\delta$, the generalization error satisfies:
$(j^{M}_{X}-P^{M}_{X})^{2}\leq 2C\sqrt{\frac{f(d)+\ln(4/\delta)}{N}}+\frac{2\ln(4/\delta)}{n}.$
Here $n$ is the number of times of generating responses for training, $N$ is the size of training set, $d$ is the number of parameters of JPPN, $f(d)$ is a complexity measure (e.g., Rademacher complexity (Mohri, Rostamizadeh, and Talwalkar 2018)), and $C$ is a constant. This implies that JPPN’s prediction is getting closer to $P^{M}_{X}$ on test set as $n$ and $N$ increase. See App. C for the full proof.
Jailbreak Probability for Attack
Once the correlation of $\hat{p}_{X}^{M}$ and $X$ is effectively encoded, it can be used to guide $X$ towards higher jailbreak probability. With a well-trained JPPN, we propose the Jailbreak-Probability-based Attack (JPA) which generates the adversarial example $X^{\prime}$ by minimizing the following loss:
$\mathcal{L}_{JPA}=\mathcal{L}_{MSE}(j_{X}^{M},1),$ (6)
where $1$ is the target and represents 100% probability of $X$ jailbreaking $M$. The adversarial image $x^{\prime}_{img}$ is generated iteratively via the following equation:
@@
s.t. $\quad\|x^{\prime}_{img}-x_{img}\|_{\infty}\leq\epsilon,$ (8)
where $x^{\prime}_{img}$ is initialized as $x_{img}$, $\alpha$ is the step size and $\epsilon$ is the perturbation bound. After iterations of updates, the hidden states of $X^{\prime}$ are located close to those of high jailbreak probability, thereby enhancing the likelihood that $X^{\prime}$ will jailbreak $M$. See App. D for the algorithm of JPA.
The effectiveness of JPA comes from the nuanced jailbreak probability labels of inputs, which allow JPPN to learn what is crucial for achieving successful jailbreak in feature space. The essence of jailbreak probability is a kind of information about inputs, which brings knowledge about the inherent vulnerabilities of the victim MLLM. Furthermore, compared to methods aligning the outputs of models with harmful responses (Qi et al. 2024; Ying et al. 2024), JPA uses a small MLP to perform optimization on a simple regression task, which is faster and much easier to converge.
Additionally, to further boost the jailbreak ability, we use an LLM to iteratively refine $x_{txt}$ with a refinement prompt $P_{r}$ (See App. F for details). $P_{r}$ contains knowledge about how to rephrase $x_{txt}$ to make it more possible to conduct jailbreaks. We enforce a monotonicity constraint on the refinement: we test the jailbreak probability of the refined text after each iteration; if the jailbreak probability is reduced, we roll back the refinement to guarantee that the jailbreak probability only increases during rephrasing. The monotonic refinement stops when reaching the maximum number of iterations. We refer to the image-only attack as JPA and the multimodal attack as MJPA in the rest of the paper. The illustration of (M)JPA is shown in Fig. 2. See App. E for the algorithm of MJPA.
Jailbreak Probability for Defense
Similar to JPA, we can also drive $j_{X}^{M}$ closer to 0 to defend against potential jailbreaks by minimizing:
@@
where 0 represents 0% probability of $X$ jailbreaking $M$. We propose Jailbreak-Probability-based Finetuning (JPF), which minimizes Eq. 9 on MLLM parameters . JPF updates the MLLM parameters via:
@@
where $\theta$ is the parameters of $M$ and $\beta$ is the learning rate. JPF is illustrated in Fig. 2.
Experiments
Experimental Setup
@@
Datasets. Zhou et al. (Zhou et al. 2024b) introduce a dataset which consists of three datasets, including AdvBench (Zou et al. 2023), StrongREJECT (Souly et al. 2024) and JailbreakBench (Chao et al. 2024). We refer to this dataset as Zhou’s dataset. Zhou’s dataset is randomly split into train set and test set with ratios of 0.8 and 0.2. For all texts in the Zhou’s dataset, we combine them with a photo of a panda to form image-text pairs, which is the same as VAP (Qi et al. 2024) and BAP (Ying et al. 2024). We also use MM-SafetyBench (Liu et al. 2024a), SafeBench (Gong et al. 2023) and VLGuard (Zong et al. 2024) in our experiments.
@@
Models. We adopt MiniGPT-4 (Vicuna 13B) (Zhu et al. 2023a), InstructBlip (Vicuna 13B) (Dai et al. 2023b), Qwen-VL-Chat (Qwen-7B) (Bai et al. 2023) InternLM-XComposer-VL (InternLM-Chat-7B) (Zhang et al. 2023), DeepSeek-VL-1.3B-Chat (DeepSeek-LLM-1.3B) (Lu et al. 2024), GPT-4V (OpenAI 2023), GPT-4o (Hurst et al. 2024) and Qwen-VL-Plus (Bai et al. 2023) in our experiments.
@@
Criterion. We use GPT-4 (OpenAI 2024) as the judge model, which responses with ”safe” or ”unsafe” to the model outputs. See App. B for the judge prompt. We generate answers for 5 times in all experiments to reduce randomness.
Training of JPPNs
@@
Training Settings. We choose MiniGPT-4 as the MLLM to implement JPPN. JPPN is implemented as a three-layer MLP. According to the dimensions of MiniGPT-4’s hidden states, JPPN’s input dimension is set to 5120. Each Transformer block is connected to a distinct JPPN to process its respective hidden states. See App. A for JPPN architecture. We train JPPNs on Zhou’s train set. To label malicious questions, we generate multiple MiniGPT-4 responses and calculate the approximated jailbreak probability using Eq. 3 based on GPT-4 with the judge prompt in App. B. Since jailbreak probability is derived from the ratio of successful jailbreak responses, its stability depends on the number of responses. To assess the impact of the response number on training, we experiment with 5, 20, and 40 responses per input, training JPPNs-5r, JPPNs-20r, and JPPNs-40r, respectively, over 150 epochs. We use Adam (Kingma 2014) with an initial learning rate of 0.001 as the optimizer, and reduce the learning rate to 20% every 50 epochs. In practice, generating 20 examples for each input in the training set costs about 4 hours on 8 NVIDIA A100 GPUs.
Analysis on JPPN’s Accuracy. After training, we evaluate JPPN’s accuracy on Zhou’s test set. Given the inherent randomness in the value of approximated jailbreak probability, we introduce a flexible metric for evaluating accuracy, which is referred to as $\tau$-thresholded accuracy ($acc_{\tau}$). $acc_{\tau}$ is calculated with an indicator function:
$acc_{\tau}=\frac{1}{\left|D\right|}\sum_{X\in D}\mathbf{1}(X),\,\,if\left|j_{X}-\hat{p}_{X}\right|\leq\tau$ (11)